# Avoiding Catastrophe in Online Learning by Asking for Help

### v4

@@ Abstract
Most learning algorithms with formal regret guarantees assume that all mistakes are recoverable and essentially rely on trying all possible behaviors. This approach is problematic when some mistakes are catastrophic, i.e., irreparable. We propose an online learning problem where the goal is to minimize the chance of catastrophe. Specifically, we assume that the payoff in each round represents the chance of avoiding catastrophe that round and try to maximize the product of payoffs (the overall chance of avoiding catastrophe) while allowing a limited number of queries to a mentor. We first show that in general, any algorithm either constantly queries the mentor or is nearly guaranteed to cause catastrophe. However, in settings where the mentor policy class is learnable in the standard online model, we provide an algorithm whose regret and rate of querying the mentor both approach 0 as the time horizon grows. Conceptually, if a policy class is learnable in the absence of catastrophic risk, it is learnable in the presence of catastrophic risk if the agent can ask for help.
@@ 1 Introduction
There has been mounting concern over catastrophic risk from AI, including but not limited to autonomous weapon accidents (Abaimov & Martellini, 2020), bioterrorism (Mouton et al., 2024), and cyberattacks on critical infrastructure (Guembe et al., 2022). See Critch & Russell (2023) and Hendrycks et al. (2023) for taxonomies of societal-scale AI risks. In this paper, we use “catastrophe” to refer to any kind of irreparable harm. This definition also covers smaller-scale (yet still unacceptable) incidents such as serious medical errors (Rajpurkar et al., 2022), crashing a robotic vehicle (Kohli & Chadha, 2020), and discriminatory sentencing (Villasenor & Foggo, 2020).
@@
More precisely, trial-and-error-style algorithms only work when catastrophe is assumed to be impossible. This assumption manifests differently in different subtypes of online learning. In the standard online learning model, the agent’s actions have no permanent effect on the environment. [^1] More precisely, the input can depend on the agent’s previous actions, but the agent’s performance is always evaluated relative to the optimal policy’s performance on the same sequence of inputs. Online reinforcement learning allows the agent’s actions to permanently affect the environment, but typically assumes that either no action has irreversible effects (e.g., Jaksch et al. (2010)) or that the agent is reset at the start of each “episode” (e.g., Azar et al. (2017)). One could train an agent entirely in a controlled lab setting where the above assumptions do hold, but we argue that sufficiently general agents will inevitably encounter novel scenarios when deployed in the real world. Machine learning models often behave unpredictably in unfamiliar environments (see, e.g., Quionero-Candela et al. (2009)), and we do not want AI biologists or robotic vehicles to behave unpredictably.
The goal of this paper is to understand the conditions under which it is possible to formally guarantee avoidance of catastrophe in online learning. Certainly some conditions are necessary, because if the agent can only learn by trying actions directly, the problem is hopeless: any untried action could lead to paradise or disaster and the agent has no way to predict which. In the real world, however, one needn’t learn through pure trial-and-error: one can also ask for help. We think it is critical for high-stakes AI applications to employ a designated supervisor who can be asked for help. Examples include a human doctor supervising AI doctors, a robotic vehicle with a human driver who can take over in emergencies, autonomous weapons with a human operator, and many more. We hope that our work constitutes a small step in the direction of practical safety guarantees for such applications.
@@ 1.1 Our model
We propose an online learning model of avoiding catastrophe with mentor help. On each time step, the agent observes an input, selects an action (or queries the mentor), and obtains a payoff. Each payoff represents the probability of avoiding catastrophe on that time step (conditioned on no prior catastrophe). The agent’s goal is to maximize the product of payoffs, which is equal to the overall probability of avoiding catastrophe. [^2] Conditioning on no prior catastrophe means we do not need to assume that these probabilities are independent (and if catastrophe has already occurred, this time step does not matter). This is due to the chain rule of probability. As is standard in online learning, we consider the product of payoffs obtained while learning, not the product of payoffs of some final policy.
The (possibly suboptimal) mentor has a fixed policy, and when queried, the mentor illustrates their policy’s action for the current input. We desire an agent whose regret – defined as the gap between the mentor’s performance and the agent’s performance – approaches zero as the time horizon $T$ grows. In other words, with enough time, the agent should avoid catastrophe nearly as well as the mentor. We also expect the agent to become self-sufficient over time: formally, the number of queries to the mentor should be sublinear in $T$, or equivalently, the rate of querying the mentor should go to zero.
@@ 1.2 Our assumptions
The agent needs some way to make inferences about unqueried inputs in order to decide when to ask for help. Much past work has used Bayesian inference, which suffers tractability issues in complex environments. [^3] For the curious reader, Betancourt (2018) provides a thorough treatment. See also Section 2. We instead assume that the mentor policy satisfies what we call local generalization: informally, if the mentor told us that an action was safe for a similar input, then that action is probably also safe for the current input (see Section 3 for a formal definition and further discussion). This captures the intuition that one can transfer knowledge between similar situations. Unlike Bayesian inference, local generalization only requires computing distances and is compatible with any input space which admits a distance metric.
Unlike the standard online learning model, we assume that the agent does not observe payoffs. This is because the payoff in our model represents the chance of avoiding catastrophe on that time step. In the real world, one only observes whether catastrophe occurred, not its probability. [^4] One may be able to detect “close calls” in some cases, but observing the precise probability seems unrealistic. One consequence of this is that the payoff function does not affect the behavior of the algorithm, so our positive results hold simultaneously for all payoff functions which satisfy local generalization.
@@ 1.3 Standard online learning
An overview of standard online learning is in order before discussing our results. In the standard model, the agent observes an input on each time step and must choose an action. An adversary then reveals the correct action, which results in some payoff to the agent. The goal is sublinear regret with respect to the sum of payoffs, or equivalently, the average regret per time step should go to 0 as $T\to\infty$. Figure 1 delineates the precise differences between the standard model and our model.
If the adversary’s choices are unconstrained, the problem is hopeless: if the adversary determines the correct action on each time step randomly and independently, the agent can do no better than random guessing. However, sublinear regret becomes possible if (1) the hypothesis class has finite Littlestone dimension (Littlestone, 1988), or (2) the hypothesis class has finite VC dimension (Vapnik & Chervonenkis, 1971) and the input is $\sigma$-smooth [^5] Informally, the adversary chooses a distribution over inputs instead of a precise input. See Section 3. (Haghtalab et al., 2024).
The goal of sublinear regret in online learning implicitly assumes catastrophe is impossible: the agent can make arbitrarily many (and arbitrarily costly) mistakes as long as the average regret per time step goes to 0. In contrast, we demand subconstant regret: the total probability of catastrophe should go to 0. Furthermore, standard online learning allows the agent to observe payoffs on every time step, while our agent only receives feedback on time steps with queries. However, access to a mentor (and local generalization) allows our agent to learn without trying actions directly, which is enough to offset all of the above disadvantages.
Objective Regret goal Feedback Mentor Local generalization
Standard model Sum of payoffs Sublinear Every time step No No
Our model Product of payoffs Subconstant Only from queries Yes Yes
Figure 1: Comparison between the standard online learning model and our model.
@@ 1.4 Our results
More precisely, we first show that in general, any algorithm with sublinear queries to the mentor has arbitrarily poor regret in the worst-case (Theorem 4.1). This means that even when the mentor can avoid catastrophe with certainty, any algorithm either needs excessive supervision or is nearly guaranteed to cause catastrophe. Unlike online learning where the general impossibility result is trivial (the agent might as well guess randomly given an unconstrained adversary), local generalization significantly limits the adversary’s power and necessitates a careful analysis.
Next, we present a simple algorithm whose total regret and rate of querying the mentor both go to 0 as $T\to\infty$ when either (1) the mentor policy class has finite Littlestone dimension or (2) the mentor policy class has finite VC dimension and the input sequence is $\sigma$-smooth. Our algorithm can handle a unbounded input space and does not need detailed access to the feature embedding, instead relying on two simple operations. It does need to know the mentor policy class, as is standard in online learning. We initially prove the theorem for binary actions (Theorem 5.2) and then reduce learning with many actions to the binary action case (Theorem C.1).
Along the way, we prove that the same subconstant bound holds for standard additive regret (Theorem 5.3). Essentially, our techniques are equally effective for maximizing the sum of payoffs and the product of payoffs. We emphasize the multiplicative objective due to our motivation of avoiding catastrophe, but our subconstant additive regret bound may also be of interest. In summary, the combination of a mentor and local generalization allows us to reduce the regret by an entire factor of $T$, resulting in subconstant regret (multiplicative or additive) instead of the typical sublinear regret.
@@ 2 Related work
Learning with irreversible costs.
Despite the ubiquity of irreparable/irreversible costs in the real world, theoretical work on this topic remains limited. This may be due to the fundamental modeling question of how the agent should learn about novel inputs or actions without actually trying them.
@@
Bayesian inference combined with asking for help is studied by Cohen et al. (2021); Cohen & Hutter (2020); Kosoy (2019); Mindermann et al. (2018). It is also worth mentioning Hadfield-Menell et al. (2017); Moldovan & Abbeel (2012); Turchetta et al. (2016), who utilize Bayesian inference in the context of safe (online) reinforcement learning without asking for help (and without regret bounds).
We are only aware of two papers which theoretically address irreversibility without Bayesian inference: Grinsztajn et al. (2021) and Maillard et al. (2019). The former proposes to sample trajectories and learn reversibility based on temporal consistency between states: intuitively, if $s_{1}$ always precedes $s_{2}$, we can infer that $s_{1}$ is unreachable from $s_{2}$. Although the paper theoretically grounds this intuition, there is no formal regret guarantee. The latter presents an algorithm which asks for help in the form of rollouts from the current state. However, the regret bound and number of rollouts are both linear in the worst case, due to the dependence on the $\gamma^{*}$ parameter which roughly captures how bad an irreversible action can be. In contrast, our algorithm achieves good regret even when actions are maximally bad.
@@
CMDPs (Altman, 2021) require the agent to maximize reward while also satisfying safety constraints. See Gu et al. (2023); Zhao et al. (2023); Wachi et al. (2024) for surveys. The two most relevant papers are Liu et al. (2021) and Stradi et al. (2024), both of which provide algorithms guaranteed to satisfy initially unknown safety constraints. Since neither paper allows external help, they require strong assumptions to make the problem tractable: the aforementioned results assume that the agent (1) knows a strictly safe policy upfront (i.e., a policy which satisfies the safety constraints with slack), (2) is periodically reset, and (3) observes the safety costs. In contrast, our agent has no prior knowledge, is never reset, and never observes payoffs.
Online learning.
@@
See Cesa-Bianchi & Lugosi (2006) and Chapter 21 of Shalev-Shwartz & Ben-David (2014) for introductions to online learning. A classical result states that sublinear regret is possible iff the hypothesis class has finite Littlestone dimension (Littlestone, 1988). However, even some simple hypothesis classes have infinite Littlestone dimension, such as the class of thresholds on $[0,1]$ (Example 21.4 in Shalev-Shwartz & Ben-David (2014)). Recently, Haghtalab et al. (2024) showed that if the adversary only chooses a distribution over inputs rather than the precise input, only finite VC dimension (Vapnik & Chervonenkis, 1971) is needed for sublinear regret. Specifically, they assume that each input is sampled from a distribution whose concentration is upper bounded by $\frac{1}{\sigma}$ times the uniform distribution. This framework is known as smoothed analysis, originally due to Spielman & Teng (2004).
Multiplicative objectives.
@@
Although online learning traditionally studies the sum of payoffs, there is some work on maximizing the product of payoffs, or equivalently the sum of logarithms (Chapter 9 of Cesa-Bianchi et al. (2017)). However, these regret bounds are still sublinear in $T$, in comparison to our subconstant regret bounds. Also, that work still assumes that payoffs are observed on every time step, while our agent only receives feedback in response to queries (Figure 1).
Barman et al. (2023) recently provided regret bounds for a multiplicative objective in a multi-armed bandit problem, but their objective is the geometric mean of payoffs instead of the product. Interpreted in our context, their regret bounds imply that the average chance of catastrophe goes to zero, while we guarantee that the total chance of catastrophe goes to zero. This distinction is closely related to the difference between subconstant and sublinear regret discussed in Section 1.3.
Active learning and imitation learning.
@@
Our assumption that the agent only receives feedback in response to queries falls under the umbrella of active learning (Hanneke, 2014). This contrasts with passive learning, where the agent receives feedback automatically. Although ideas from active learning could be useful in our domain, we are not aware of any results from that literature which account for irreversible costs. The process of the agent learning from a mentor is also reminiscent of imitation learning (Osa et al., 2018), but we are not aware of any relevant technical implications.
@@ 3 Model
Let $\mathbb{N}$ refer to the set of strictly positive integers and let $T\in\mathbb{N}$ be the time horizon. Let $\boldsymbol{x}=(x_{1},x_{2},\dots x_{T})\in\mathcal{X}^{T}$ be the sequence of inputs. In the adversarial setting, each $x_{t}$ can have arbitrary dependence on the events of prior time steps. In the smoothed setting, the adversary only chooses the distribution $x_{t}$ from which $x_{t}$ is sampled. Formally, a distribution $\mathcal{D}$ over $\mathcal{X}$ is $\sigma$-smooth if for any $X\subseteq\mathcal{X}$, $\mathcal{D}(X)\leq\frac{1}{\sigma}U(X)$. (In the smoothed setting, we assume that $\mathcal{X}$ supports a uniform distribution $U$ [^6] For example, $\mathcal{X}$ having finite Lebesgue measure is sufficient. Note that this does not imply boundedness. Alternatively, $\sigma$-smoothness can be defined with respect to a different distribution, as long as the Radon-Nikodym derivative is uniformly bounded; see Definition 1 of Block et al. (2022).) If each $x_{t}$ is sampled from a $\sigma$-smooth $\mathcal{D}_{t}$, we say that $\boldsymbol{x}$ is $\sigma$-smooth. The sequence $\boldsymbol{\mathcal{D}}=\mathcal{D}_{1},\dots,\mathcal{D}_{T}$ can still be adaptive, i.e., the choice of $\mathcal{D}_{t}$ can depend on the events of prior time steps.
Actions and payoffs.
@@
Let $\mathcal{Y}$ be a finite set of actions. There also exists a special action $\tilde{y}$ which corresponds to querying the mentor. For $k\in\mathbb{N}$, let $[k]=\{1,2,\dots,k\}$. On each time step $t\in[T]$, the agent must select action $y_{t}\in\ \mathcal{Y}\cup\{\tilde{y}\}$, which generates a payoff. Let $\boldsymbol{y}=(y_{1},\dots,y_{T})$. We allow the payoff function to vary between time steps: let $\boldsymbol{\mu}=(\mu_{1},\dots,\mu_{T})\in(\mathcal{X}\times\mathcal{Y}\to[0,1])^{T}$ be the sequence of payoff functions. [^7] Variation in $\mu_{t}$ is not a core part of our problem, but we allow it because the exact same proof holds. Since the mentor policy is fixed, local generalization limits how much $\mu_{t}$ can vary between time steps. Then $\mu_{t}(x_{t},y_{t})\in[0,1]$ is the agent’s payoff at time $t$. Like $\boldsymbol{\mathcal{D}}$, we allow $\boldsymbol{\mu}$ to be adaptive. Unless otherwise noted, all expectations are over the agent’s randomization (if any) and the randomization in $\boldsymbol{x}$ (if any).
Asking for help.
@@
The mentor is endowed with a (not necessarily optimal) policy $\pi^{m}:\mathcal{X}\to\mathcal{Y}$. When action $\tilde{y}$ is chosen, the mentor informs the agent of the action $\pi^{m}(x_{t})$ and the agent obtains payoff $\mu_{t}(x_{t},\pi^{m}(x_{t}))$. For brevity, let $\mu_{t}^{m}(x)=\mu_{t}(x,\pi^{m}(x))$. The agent never observes payoffs: the only way to learn about $\boldsymbol{\mu}$ is by querying the mentor.
@@
We would like an algorithm which becomes “self-sufficient” over time: the rate of querying the mentor should go to 0 as $T\to\infty$, or equivalently, the cumulative number of queries should be sublinear in $T$. Formally, let $Q_{T}(\boldsymbol{\mu},\pi^{m})=\{t\in[T]:y_{t}=\tilde{y}\}$ be the random variable denoting the set of time steps with queries. Then we say that the (expected) number of queries is sublinear in $T$ if $\sup_{\boldsymbol{\mu},\pi^{m}}\operatorname*{\mathbb{E}}[|Q_{T}(\boldsymbol{\mu},\pi^{m})|]\in o(T)$. In other words, there must exist $g:\mathbb{N}\to\mathbb{N}$ such that $g(T)\in o(T)$ and $\sup_{\boldsymbol{\mu},\pi^{m}}\operatorname*{\mathbb{E}}[|Q_{T}(\boldsymbol{\mu},\pi^{m})|]\leq g(T)$. [^8] One could instead consider the worst-case number of queries, but this distinction does not affect whether subconstant regret is achievable (Proposition E.2). For brevity, we will usually write $Q_{T}=Q_{T}(\boldsymbol{\mu},\pi^{m})$.
Local generalization.
@@
We assume that the mentor policy permits “local generalization”. Informally, if the agent is given an input $x$, taking the mentor action for a similar input $x^{\prime}$ is almost as good as taking the mentor action for $x$. Formally, we assume $\mathcal{X}\subseteq\mathbb{R}^{n}$ and there exists $L>0$ such that for all $x,x^{\prime}\in\mathcal{X}$ and $t\in[T]$, $|\mu_{t}^{m}(x)-\mu_{t}(x,\pi^{m}(x^{\prime}))|\leq L||x-x^{\prime}||$, where |||||$\cdot$||| | ⋅ | | denotes Euclidean distance. [^9] One could also allow a generic metric space; our assumption of $\mathcal{X}\subseteq\mathbb{R}^{n}$ is only for convenience. This represents the ability to transfer knowledge between similar inputs:

### v5

@@ Abstract
Most learning algorithms with formal regret guarantees assume that all mistakes are recoverable and essentially rely on trying all possible behaviors. This approach is problematic when some mistakes are catastrophic, i.e., irreparable. We propose an online learning problem where the goal is to minimize the chance of catastrophe. Specifically, we assume that the payoff in each round represents the chance of avoiding catastrophe in that round and try to maximize the product of payoffs (the overall chance of avoiding catastrophe) while allowing a limited number of queries to a mentor. We also assume that the agent can transfer knowledge between similar inputs. We first show that in general, any algorithm either queries the mentor at a linear rate or is nearly guaranteed to cause catastrophe. However, in settings where the mentor policy class is learnable in the standard online model, we provide an algorithm whose regret and rate of querying the mentor both approach 0 as the time horizon grows. Although our focus is the product of payoffs, we provide matching bounds for the typical additive regret. Conceptually, if a policy class is learnable in the absence of catastrophic risk, it is learnable in the presence of catastrophic risk if the agent can ask for help.
online learning, AI safety, asking for help, irreversibility
@@ 1 Introduction
There has been mounting concern over catastrophic risk from AI, including but not limited to autonomous weapon accidents (Abaimov & Martellini, 2020), bioterrorism (Mouton et al., 2024), cyberattacks on critical infrastructure (Guembe et al., 2022), and loss of control (Bengio et al., 2024). See Critch & Russell (2023) and Hendrycks et al. (2023) for taxonomies of societal-scale AI risks. In this paper, we use “catastrophe” to refer to any kind of irreparable harm. In addition to the large-scale risks above, our definition also covers smaller-scale (yet still unacceptable) incidents such as serious medical errors (Rajpurkar et al., 2022), crashing a robotic vehicle (Kohli & Chadha, 2020), and discriminatory sentencing (Villasenor & Foggo, 2020).
@@
More precisely, trial-and-error-style algorithms only work when catastrophe is assumed to be impossible. This assumption can take multiple forms, such as that the agent’s actions do not affect future inputs (e.g., Slivkins, 2011), that no action has irreversible effects (e.g., Jaksch et al., 2010) or that the environment is reset at the start of each “episode” (e.g., Azar et al., 2017). One could train an agent entirely in a controlled lab setting where one of those assumptions does hold, but we argue that sufficiently general agents will inevitably encounter novel scenarios when deployed in the real world. Machine learning models often behave unpredictably in unfamiliar environments (see, e.g., Quiñonero-Candela et al., 2022), and we do not want AI biologists or robotic vehicles to behave unpredictably.
The goal of this paper is to understand the conditions under which it is possible to formally guarantee avoidance of catastrophe in online learning. Certainly some conditions are necessary, because the problem is hopeless if the agent must rely purely on trial-and-error: any untried action could lead to paradise or disaster and the agent has no way to predict which. In the real world, however, one need not learn through pure trial-and-error: one can also ask for help. We think it is critical for high-stakes AI applications to employ a designated supervisor who can be asked for help. Examples include a human doctor supervising AI doctors, a robotic vehicle with a human driver who can take over in emergencies, autonomous weapons with a human operator, and many more. We hope that our work constitutes a step in the direction of practical safety guarantees for such applications.
@@ 1.1 Our model
We propose an online learning model of avoiding catastrophe with mentor help. On each time step, the agent observes an input, selects an action or queries the mentor, and obtains a payoff. Each payoff represents the probability of avoiding catastrophe on that time step (conditioned on no prior catastrophe). The agent’s goal is to maximize the product of payoffs, which is equal to the overall probability of avoiding catastrophe. [^1] Conditioning on no prior catastrophe means we do not need to assume that these probabilities are independent (and if catastrophe has already occurred, this time step does not matter). This is due to the chain rule of probability. As is standard in online learning, we consider the product of payoffs obtained while learning, not the product of payoffs of some final policy.
The (possibly suboptimal) mentor has a stationary policy, and when queried, the mentor illustrates their policy’s action for the current input. We want the agent’s regret – defined as the gap between the mentor’s performance and the agent’s performance – to go to zero as the time horizon $T$ grows. In other words, with enough time, the agent should avoid catastrophe nearly as well as the mentor. We also expect the agent to become self-sufficient over time: the number of queries to the mentor should be sublinear in $T$, or equivalently, the rate of querying the mentor should go to zero.
@@ 1.2 Our assumptions
The agent needs some way to make inferences about unqueried inputs in order to decide when to ask for help. Much past work has used Bayesian inference, which suffers from tractability issues in complex environments. [^2] For the curious reader, Betancourt (2018) provides a thorough treatment. See also Section 2. We instead assume that the mentor policy satisfies a novel property that we call local generalization: informally, if the mentor told us that an action was safe for a similar input, then that action is probably also safe for the current input. For example, if it is safe to ignore a 3 mm spot on an X-ray, it is likely (but not certainly) also safe to ignore a 3.1 mm spot with the same density, location, etc. Unlike Bayesian inference, local generalization only requires computing distances and is compatible with any input space that admits a distance metric. See Section 5.2 for further discussion of local generalization.
Unlike the standard online learning model, we assume that the agent does not observe payoffs. This is because the payoff in our model represents the chance of avoiding catastrophe on that time step. In the real world, one only observes whether catastrophe occurred, not its probability. [^3] One may be able to detect “close calls” in some cases, but observing the precise probability seems unrealistic.
@@ 1.3 Standard online learning
To properly understand our results, it is important to understand standard online learning. In the standard model, the agent observes an input on each time step and must choose an action. An adversary then reveals the correct action, which results in some payoff to the agent. The goal is sublinear regret with respect to the sum of payoffs, or equivalently, the average regret per time step should go to 0 as $T\to\infty$. Table 1 delineates the precise differences between the standard model and our model.
If the adversary’s choices are unconstrained, the problem is hopeless: if the adversary determines the correct action on each time step randomly and independently, the agent can do no better than random guessing. However, sublinear regret becomes possible if (1) the hypothesis class has finite Littlestone dimension (Littlestone, 1988), or (2) the hypothesis class has finite VC dimension (Vapnik & Chervonenkis, 1971) and the input is $\sigma$-smooth [^4] Informally, the adversary chooses a distribution over inputs instead of a precise input. See Section 3 for the formal definition. (Haghtalab et al., 2024).
The goal of sublinear regret in online learning implicitly assumes catastrophe is impossible: the agent can make arbitrarily many (and arbitrarily costly) mistakes as long as the average regret per time step goes to 0. In contrast, we demand subconstant regret: the total probability of catastrophe should go to 0. Furthermore, standard online learning allows the agent to observe payoffs on every time step, while our agent only receives feedback on time steps with queries. However, the combination of a mentor and local generalization allows our agent to learn without trying actions directly, which is enough to offset all of the above disadvantages.
Table 1: Comparison between the standard online learning model and our model.
Standard model Our model
Objective Sum of payoffs Product of payoffs
Regret goal Sublinear Subconstant
Feedback Every time step Only from queries
Mentor No Yes
Local generalization No Yes
@@ 1.4 Our results
We first show that in general, any algorithm with sublinear queries to the mentor has unbounded regret in the worst-case (Theorem 4.1). As a corollary, even when the mentor can avoid catastrophe with certainty, any algorithm either needs extensive supervision or is nearly guaranteed to cause catastrophe (Corollary 4.1.1).
Our primary result is a simple algorithm whose total regret and rate of querying the mentor both go to 0 as $T\to\infty$ when either (1) the mentor policy class has finite Littlestone dimension or (2) the mentor policy class has finite VC dimension and the input sequence is $\sigma$-smooth (Theorem 5.2). Conceptually, the algorithm has two components: (1) for “in-distribution” inputs, run a standard online learning algorithm (adjusted to account for only receiving feedback in response to queries), and (2) for “out-of-distribution” inputs, ask for help. Our algorithm can handle an unbounded input space and does not need to know the local generalization constant.
Although we focus on the product of payoffs, we show that the results above (both positive and negative) hold for the typical additive regret as well. In fact, we show that multiplicative regret and additive regret are tightly related in our setting (Lemma A.1).
In summary, the combination of local generalization and a mentor allows us to reduce the regret by an entire factor of $T$, resulting in subconstant regret (multiplicative or additive) instead of the typical sublinear regret.
@@ 2 Related work
Learning with irreversible costs. Despite the ubiquity of irreversible costs in the real world, theoretical work on this topic remains limited. This may be due to the fundamental modeling question of how the agent should learn about novel inputs or actions without trying them directly.
@@
Bayesian inference combined with asking for help is studied by Cohen et al. (2021); Cohen & Hutter (2020); Kosoy (2019); Mindermann et al. (2018). We also mention Hadfield-Menell et al. (2017); Moldovan & Abbeel (2012); Turchetta et al. (2016), who utilize Bayesian inference in the context of safe (online) reinforcement learning without asking for help (and without regret bounds).
We are only aware of two papers that theoretically address irreversibility without Bayesian inference: Grinsztajn et al. (2021) and Maillard et al. (2019). The former proposes to sample trajectories and learn reversibility based on temporal consistency between states: intuitively, if $s_{1}$ always precedes $s_{2}$, we can infer that $s_{1}$ is unreachable from $s_{2}$. Although the paper theoretically grounds this intuition, there is no formal regret guarantee. The latter presents an algorithm which asks for help in the form of rollouts from the current state. However, the regret bound and number of rollouts are both linear in the worst case, due to the dependence on the $\gamma^{*}$ parameter which roughly captures how bad an irreversible action can be. In contrast, our algorithm achieves good regret even when actions are maximally bad.
@@
Safe reinforcement learning (RL). The safe RL problem is typically formulated as a constrained Markov Decision Process (CMDP) (Altman, 2021). In CMDPs, the agent must maximize reward while also satisfying safety constraints. See Gu et al. (2024); Zhao et al. (2023); Wachi et al. (2024) for surveys. The two most relevant safe RL papers are Liu et al. (2021) and Stradi et al. (2024), both of which provide algorithms guaranteed to satisfy initially unknown safety constraints. Since neither paper allows external help, they require strong assumptions to make the problem tractable: the aforementioned results assume that the agent (1) knows a strictly safe policy upfront (i.e., a policy which satisfies the safety constraints with slack), (2) is periodically reset, and (3) observes the safety costs. In contrast, our agent has no prior knowledge, is never reset, and never observes payoffs.
@@
Online learning. See Cesa-Bianchi & Lugosi (2006) and Chapter 21 of Shalev-Shwartz & Ben-David (2014) for introductions to online learning. A classical result states that sublinear regret is possible if and only if the hypothesis class has finite Littlestone dimension (Littlestone, 1988). However, even some simple hypothesis classes have infinite Littlestone dimension, such as the class of thresholds on $[0,1]$ (Example 21.4 in Shalev-Shwartz & Ben-David, 2014). Recently, Haghtalab et al. (2024) showed that if the adversary only chooses a distribution over inputs rather than the precise input, only finite VC dimension (Vapnik & Chervonenkis, 1971) is needed for sublinear regret. Specifically, they assume that each input is sampled from a distribution whose concentration is upper bounded by $\frac{1}{\sigma}$ times the uniform distribution. This framework is known as smoothed analysis, originally due to Spielman & Teng (2004).
@@
Multiplicative objectives. Although online learning traditionally studies the sum of payoffs, there is some work which aims to maximize the product of payoffs (or equivalently, the sum of logarithms). See, e.g., Chapter 9 of Cesa-Bianchi & Lugosi (2006). However, these regret bounds are still sublinear in $T$, in comparison to our subconstant regret bounds. Also, like most online learning work, those results assume that payoffs are observed on every time step. In contrast, our agent only receives feedback in response to queries (Table 1) and never observes payoffs. Barman et al. studied a multiplicative objective in a multi-armed bandit context, but their objective is the geometric mean of payoffs instead of the product. Interpreted in our context, their regret bounds imply that the average chance of catastrophe goes to zero, while we guarantee that the total chance of catastrophe goes to zero. This distinction is closely related to the difference between subconstant and sublinear regret.
@@
Active learning and imitation learning. Our assumption that the agent only receives feedback in response to queries falls under the umbrella of active learning (Hanneke, 2014). This contrasts with passive learning, where the agent receives feedback automatically. The way our agent learns from the mentor is also reminiscent of imitation learning (Osa et al., 2018). Although ideas from these areas could be useful in our setting, we are not aware of any results from that literature which account for irreversible costs.
@@ 3 Model
Inputs. Let $\mathbb{N}$ denote the strictly positive integers and let $T\in\mathbb{N}$ be the time horizon. Let $\boldsymbol{x}=(x_{1},x_{2},\dots,x_{T})\in\mathcal{X}^{T}$ be the sequence of inputs. In the fully adversarial setting, each $x_{t}$ can have arbitrary (possibly randomized) dependence on the events of prior time steps. In the smoothed setting, the adversary only chooses the distribution $\mathcal{D}_{t}$ from which $x_{t}$ is sampled. Formally, a distribution $\mathcal{D}$ over $\mathcal{X}$ is $\sigma$-smooth if for any $S\subseteq\mathcal{X}$, $\mathcal{D}(S)\leq\frac{1}{\sigma}U(S)$. (In the smoothed setting, we assume that $\mathcal{X}$ supports a uniform distribution $U$ [^5] For example, it suffices for $\mathcal{X}$ to have finite Lebesgue measure. Note that this does not imply boundedness. Alternatively, $\sigma$-smoothness can be defined with respect to a different distribution; see Definition 1 of Block et al. (2022).) If each $x_{t}$ is sampled from a $\sigma$-smooth $\mathcal{D}_{t}$, we say that $\boldsymbol{x}$ is $\sigma$-smooth. The sequence $\boldsymbol{\mathcal{D}}=\mathcal{D}_{1},\dots,\mathcal{D}_{T}$ can still be adaptive, i.e., the choice of $\mathcal{D}_{t}$ can depend on the events of prior time steps.
@@
Actions and payoffs. Let $\mathcal{Y}$ be a finite set of actions. There also exists a special action $\tilde{y}$ which corresponds to querying the mentor. For $k\in\mathbb{N}$, let $[k]=\{1,2,\dots,k\}$. On each time step $t\in[T]$, the agent must select action $y_{t}\in\ \mathcal{Y}\cup\{\tilde{y}\}$, which generates a payoff. Let $\boldsymbol{y}=(y_{1},\dots,y_{T})$. We allow the payoff function to vary between time steps: let $\boldsymbol{\mu}=(\mu_{1},\dots,\mu_{T})\in(\mathcal{X}\times\mathcal{Y}\to[0,1])^{T}$ be the sequence of payoff functions. Then $\mu_{t}(x_{t},y_{t})\in[0,1]$ is the agent’s payoff at time $t$. Like $\boldsymbol{\mathcal{D}}$, we allow $\boldsymbol{\mu}$ to be adaptive. Unless otherwise noted, all expectations are over any randomization in the agent’s decisions, any randomization in $\boldsymbol{x}$, and any randomization in the adaptive choice of $\boldsymbol{\mu}$.
@@
Asking for help. The mentor is endowed with a (possibly suboptimal) policy $\pi^{m}:\mathcal{X}\to\mathcal{Y}$. When action $\tilde{y}$ is chosen, the mentor informs the agent of the action $\pi^{m}(x_{t})$ and the agent obtains payoff $\mu_{t}(x_{t},\pi^{m}(x_{t}))$. For brevity, let $\mu_{t}^{m}(x)=\mu_{t}(x,\pi^{m}(x))$. The agent never observes payoffs: the only way to learn about $\boldsymbol{\mu}$ is by querying the mentor.
@@
We would like an algorithm which becomes “self-sufficient” over time: the rate of querying the mentor should go to 0 as $T\to\infty$, or equivalently, the cumulative number of queries should be sublinear in $T$. Formally, let $Q_{T}(\boldsymbol{\mu},\pi^{m})=\{t\in[T]:y_{t}=\tilde{y}\}$ be the random variable denoting the set of time steps with queries. Then we say that the (expected) number of queries is sublinear in $T$ if $\sup_{\boldsymbol{\mu},\pi^{m}}\operatorname*{\mathbb{E}}[|Q_{T}(\boldsymbol{\mu},\pi^{m})|]\in o(T)$. In other words, there must exist $g:\mathbb{N}\to\mathbb{N}$ which does not depend on $\boldsymbol{\mu}$ or $\pi^{m}$ such that $g(T)\in o(T)$ and $\sup_{\boldsymbol{\mu},\pi^{m}}\operatorname*{\mathbb{E}}[|Q_{T}(\boldsymbol{\mu},\pi^{m})|]\leq g(T)$. For brevity, we will usually write $Q_{T}=Q_{T}(\boldsymbol{\mu},\pi^{m})$.
@@
Local generalization. We assume that $\boldsymbol{\mu}$ and $\pi^{m}$ satisfy local generalization. Informally, if the agent is given an input $x$, taking the mentor action for a similar input $x^{\prime}$ is almost as good as taking the mentor action for $x$. Formally, we assume $\mathcal{X}\subseteq\mathbb{R}^{n}$ and there exists $L>0$ such that for all $x,x^{\prime}\in\mathcal{X}$ and $t\in[T]$, $|\mu_{t}^{m}(x)-\mu_{t}(x,\pi^{m}(x^{\prime}))|\leq L||x-x^{\prime}||$, where |||||$\cdot$||| | ⋅ | | denotes Euclidean distance. This represents the ability to transfer knowledge between similar inputs: